01:14 25Bias_Bino: lbcb.20080601.011506.fits - lbcb.20080601.012831.fits
lbcr.20080601.011531.fits - lbcr.20080601.011627.fits [lbcr froze after 6 frames due to lbc communication error; ran lbckill/lbcstart to solve issue, but lbcr hung during reinitialization (rotator issues). cycled power to lbcr via powercontrol panel, which appears to have restored lbcr.]

01:45 25Bias_Bino: lbcb.20080601.014641.fits - lbcb.20080601.014937.fits [stopped OB due to lbcr error!]
lbcr.20080601.014633.fits [lbcr froze again, after 1 bias frame!]
[turned lbcr power off via powercontrol panel, and rebooted lbccontrol, which did not solve problem....with Olga's help, Ben addressed the issue --- see his notes for the details.]

02:24 dome opening...

02:50 ...currently having issues with the telescope's hbs...

02:55 low clouds and contrails in the west....not photometric, but we will soldier on....

03:08 problem with hbs is proving to be more serious (it won't stay on)....

03:19 telescope is back online...

03:25 taking single (bino) test bias frame to confirm the cooperation of the lbcr. lbcb looks good. the lbcr is now exhibiting a uniform banding pattern across all 4 chips. the oscillations (bands) are at a ~few percent level. this banding is first apparent in the frame taken at 01:46 (ut time) and not seen in the frame taken at 01:11 (ut time). the pattern varies from frame to frame.

03:41 problem with banding on lbcr is not solved, but we are slewing to the EGS in order to focus (and keeping our fingers crossed in hopes that the bands disappear of their own freewill).

03:45 running focus OB...large correction found (pupils found) and sent.

03:50 rerunning focus OB....okay in red (sending large corrections)...blue is poorly fit with lots of coma (manually adding 1000 to Z8 on blue side)

03:54 rerunning focus OB....red looks good with 0.62" seeing (sdss r-band) estimate (no corrections sent). blue arrives at large corrections, which are being sent.

03:57 rerunning focus OB....blue is getting closer!

03:59 rerunning focus OB....blue is close enough. seeing of ~0.9" in V-band.

04:02 starting first EGS OB: EGS_Y_5b.

04:15 focus on red side seems to be a bit off (~1"). auto-focus script run (by John Hill) on tech chip #2 recommends a change in focus of -500 in Z4 for lbcr. pausing OB to send focus move (after exposure 3 of 5). pause does not appear to have worked, so focus was sent ~30 seconds into the 4th exposure.

04:20 lbcb is giving ~1.1" in U-band, with reasonably round stellar images.

04:25 lbcr is now giving ~0.7" seeing in Y-band (so auto-focus script moved things in the correct direction).

04:29 refocusing (twice)...small corrections in red and moderate in blue.

04:32 running EGS_Y_5a OB. guider jumped during second(?) exposure in OB. no jump evident in any of the frames though....curious.

04:57 focusing again.....rotator failure on lbcr....rerunning OB solved the issue.

05:02 running EGS_Y_5b OB. first image has seeing of ~0.6-0.7" in Y-band.

05:28 running EGS_Y_1a OB. (no focus adjustment made).

05:52 moving to Bootes and focusing...

06:01 starting priority6_PA180: b02r OB. seeing of roughly ~3pix in Y and ~4pix in U for first image of OB.

06:27 moving to next OB and focusing...

06:32 starting priority6_PA180: b04r OB. seeing of roughly 3pix in Y-band and 4.6pix in U-band in first image of OB. second image shows poorer U-band seeing (>5pix).

06:57 moving to next OB and focusing...

07:03 starting priority6_PA180: b15r OB. first image shows E-W jump/elongation. seeing is >3.5pix in Y-band and >6.5pix in U-band.

07:30 moving to next OB and focusing....

07:37 starting priority3_PA180: b11r OB.

07:56 controller issue with lbcr: no 5th image on red side. cycled power on lbc using power control window.

08:09 moving to Hercules (Herc_9) and focusing (at PA=0)...

08:19 starting Herc_9 OB: lbcb.20080601.081950.fits is elongated in NE-SW direction. lbcb.20080601.082535.fits is slightly elongated in NE-SW direction (less than previous frame though). lbcb.20080601.083115.fits is elongated like the first frame. lbcb.20080601.083657.fits is elongated like the first and third frames. lbcb.20080601.084237.fits is elongated, but more like the second frame (not as bad as frames 1,3,4).

08:55 checking the focus....

09:00 moving to Herc10 and starting OB: seeing is improving! rotator fault on red side: two blue images taken, but only one red image. cycling power on lbc using powercontrol panel.

09:23 restarting Herc10 OB:

09:59 moving to Herc11 and focusing...

10:04 starting Herc11 OB.

10:39 running Herc 11 OB again, rather than starting a new field. stopped after 3 or 4 frames as background increased: frame 1 of OB shows a jump in image (double stars, elongated images). satellite or more specifically, an alien spaceship, in frame lbcr.20080601.104532.fits (2nd image in OB). a second ufo in frame lbcr.20080601.105118.fits (3rd image in OB). stopping OB during 5th frame.

11:04 moving to skyflat position...

11:09 taking test skyflat in B/r bands....no go, though, since blue camera appears to be hung. retried OB with no success. cycling power on lbcb. mirror was in panic state. lbcb remains unresponsive....lbckill/lbcstart seems to have awakened the lbcb.

11:35 taking test skyflat in B/r bands....another lbcr failure! ~6k counts in 1-second lbcb (B-band) flat.

11:43 taking another test skyflat in B/r bands....

11:45 taking 5 B/r skyflats (Skyflat_Br_5 OB, scaling exposure time by 0.4): ~30k on red side and ~20 on blue side in first frame.

11:49 running Skyflat_test_UsY: Y-band is saturated.

11:53 running Skyflat_UsY_5 OB, scaling exposure time by 0.2: >40k on red side and ~20k on blue side in first frame.

11:58 running Skyflat_UY_5 OB, scaling exposure time by 0.15: ~30k on blue side and saturated on red side in first frame.

12:02 moving telescope to horizon.

12:09 attempting to take bias frames....25Bias_Bino OB. lbcr froze yet again! initially because it could not talk to the mirror. The sequence that caused problems was: stowed telescope, then attempted to take biases but forgot to uncheck the "use mirror" boxes. This caused one camera to go into a stop state. Normally we might fix this by turning off/on from the power control window, but the red camera would not come up.

12:43 several attempts to turn camera off/on, lbckill/lbcstart, and rebooting CMU fail to bring up LBC-red - keeps reporting RPC timeouts

13:04 giving up on LBC-red and taking 25Bias_Bino with only blue side enabled, so at least we have the blue biases.

Night summary: overall a good night. Clouds low in NW early seemed to move away, seeing was good most of the night. Instrument and telescope problems were mostly at begin and end of the night, so calibrations are a little underdone. The RPC timeout error can be extremely frustrating as there is no obvious way to know if we are going to be able to bring the red camera up, and it makes the off/on cycle very long.

My voodoo theory is that this error is a problem when the telescope is stowed: it was happening during the afternoon and after stowing at the end of the night when I wanted to take biases.

-- XiaohuiFan - 01 Jun 2008
Topic revision: r3 - 02 Jun 2008, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback